OpenStudy (anonymous):

Find the geometric mean between 18 and 2.

OpenStudy (anonymous):

geometric mean of \(a\) and \(b\) is \(\sqrt{a\times b}\)

OpenStudy (anonymous):

in your case \[\sqrt{12\times 18}\]

zepdrix (zepdrix):

* \[\Large\rm \sqrt{18\times2}\]

OpenStudy (anonymous):

thanks